OpenStudy (anonymous):

plzzzzzzz somebody could help me to find the measure of each interior angle of a regular convex do decagon and show me the equation used in this case PLZZZZ guyssssssss

OpenStudy (wolf1728):

Total of angles in a polygon = (# of sides -2)*180 Triangle = 180 square 2*180 = 360 dodecagon = n = 12 n-2 = 10 total of all angles = 10*180= 1,800 Each angle = 1,800 / 12 = 150°

OpenStudy (anonymous):

yes thanks a lot I was multiplying and I was wrong thanks

OpenStudy (wolf1728):

Glad to help out :-)

OpenStudy (anonymous):

I appreciated this very much thanksssssss

OpenStudy (wolf1728):

Yes, your question seemed as if you REALLY wanted an answer.

OpenStudy (anonymous):

I am learning geometry and it is hard for me geometry